1. Prove that the area of a parallelogram with sides A and B is |A×B|. 
1
Expert's answer
2021-01-15T05:14:11-0500

By defenition"|A×B|=ABsin\alpha,"

Area of parallelogram equals two areas of triangles "\frac 12 AB sin \alpha," i.e. "ABsin\alpha," where "\alpha" is angle befween A and B.